The senior classes at High School A and High School B planned separate trips to New York City. The senior class at High School A rented and filled 2 vans and 4 buses with 214 students. High School B rented and filled 4 vans and 5 buses with 281 students. Every van had the same number of students in it as did the buses. How many students can a van carry? How many students can a bus carry?
Answer by jorel1380(3719)

Let v be for vans , and b be for buses. Then:
2v+4b=214
4v+5b=281
So:
4v+8b=428
4v+5b=281
3b=147
b=49
v=9
Each bus carries 49 students; each van carries 9. ☺☺☺☺